# ReSem3D: Refinable 3D Spatial Constraints via Fine-Grained Semantic Grounding for Generalizable Robotic Manipulation

### v3

@@ V Discussion
Given challenges to achieve fine-grained visual grounding and maintaining real-time closed-loop control, ReSem3D achieves better performance via three strategies: (1) Refinable Two-Stage 3D Spatial Constraint Model. Semantics-driven 3D spatial constraints are generated in two stages, enabling automatic refinement and adaptation in semantically rich and sparse environments. (2) Real-Time Closed-Loop Optimization in Joint Space. Semantics-driven 3D spatial constraints are formulated as joint-space cost functions and optimized using a real-time constrained solver running over 15 Hz, enabling reactive closed-loop control in dynamic environments. (3) MLLM-driven Automated TAMP Modeling. ReSem3D autonomously determines the suitable granularity for 3D spatial constraint modeling from natural language instructions and generates cost functions for the real-time constrained optimizer. Additionally, it decomposes multi-stage tasks and generates preconditions and postconditions, enabling temporal management and dynamic backtracking during execution.
Ultimately, the stable operation of ReSem3D requires effective coordination across functional modules. To systematically identify failure causes, we conduct independent validation for each sub-module by calculating the failure rate based on Table III. Specifically, since the VFM module lacks semantic guidance, it may fail to provide task-relevant semantic grounding in 7% of cases, thereby generating irrelevant candidates. Reasoning failures in constraint extraction and refinement, constituting 16% and 28% of failures respectively, arise from insufficient cross-modal semantic grounding and limited spatial reasoning. The Constrained Optimization module often suffers from slow convergence under significant pose variation in 11% of occurrences, due to multi-objective complexity. The TAPNext module exhibits reduced tracking accuracy under occlusion or inaccurate point cloud mappings in 38% of situations, thereby breaking the perception-action loop.
Figure 9: Comparision between Copa and ReSem3D (Ours) in visual grounding.
To address these challenges, future work will explore the integration of MLLM-based self-correction strategies, global optimization algorithms, and multi-view fusion methods to further enhance the robustness and adaptability of the framework.
@@ VI Conclusion
In this work, we present ReSem3D, a unified robotic manipulation framework for semantically diverse environments. By integrating MLLMs and VFMs, the framework automatically constructs semantically grounded two-stage 3D spatial constraints. These are embedded into a hierarchical TAMP model and formulated as joint-space cost functions, enabling real-time and robust closed-loop control. The proposed approach facilitates hierarchical constraint modeling from semantically rich to sparse environments, enabling multi-stage task execution and condition-aware dynamic backtracking. Extensive simulation and real-world experiments demonstrate that ReSem3D outperforms existing baselines in zero-shot manipulation tasks across semantically diverse environments and different robotic platforms.

### v4

@@ V Discussion
Given challenges to achieve fine-grained semantic grounding and maintaining real-time closed-loop manipulation, ReSemAct achieves better performance via two strategies: (1) Semantic Structuring and Affordance Refinement. A unified semantic affordance description is introduced to guide the affordance refinement, enabling fine-grained semantic grounding and robotic manipulation. (2) MLLM-driven Automated TAMP Modeling. ReSemAct autonomously achieves fine-grained semantic grounding via SSAR under multimodal reasoning and generates cost functions for the constrained optimizer. Additionally, it decomposes multi-stage tasks and generates preconditions and postconditions, enabling temporal management and dynamic backtracking during execution.
Figure 9: Semantic grounding comparision between Copa and ReSemAct.
Figure 10: ReSemAct Ablation experiments on Four Tasks.
Ultimately, the stable operation of ReSemAct requires effective coordination across functional modules. To systematically identify failure causes, we conduct independent validation for each sub-module by calculating the failure rate based on Table III. Specifically, since the VFM module lacks semantic guidance, it may fail to provide task-relevant semantic grounding in 7% of cases, thereby generating irrelevant candidates. Reasoning failures in semantic structuring and affordance refinement, constituting 16% and 28% of failures respectively, arise from insufficient cross-modal semantic grounding and limited spatial reasoning. The Constrained Optimization module often suffers from slow convergence under significant pose variation in 11% of occurrences, due to multi-objective complexity. The TAPNext module exhibits reduced tracking accuracy under occlusion or inaccurate point cloud mappings in 38% of situations, thereby breaking the control loop. To tackle these challenges, future work will explore the integration of MLLM-based self-correction strategies, global optimization algorithms, and multi-view fusion methods to further enhance the robustness and adaptability of the framework.
@@ VI Conclusion
In this work, we present ReSemAct, a unified robotic manipulation framework for semantically diverse environments. By leveraging synergistic reasoning between MLLMs and VFMs, we introduces Semantic Structuring and Affordance Refinement (SSAR) to achieve fine-grained robot manipulation. The refined targets are embedded into a automated TAMP framework and formulated as optimizable cost functions, enabling real-time closed-loop control. The proposed approach supports multi-stage task execution and condition-aware dynamic backtracking. Extensive experiments demonstrate that ReSemAct significantly outperforms existing methods across semantically diverse environments and multiple robotic platforms.